OpenStudy (anonymous):

Solve the following system of equations by using the substitution method. x = y + 3 2x + y = 9 A. (4, 1) B. (5, 2 C. (3, -2) D. Inconsistent

OpenStudy (anonymous):

Same thing as last time, substitute one equation into the other and solve the variable.

OpenStudy (anonymous):

x-intercept = 9/2 = 4.5 y-intercept = 9/1 = 9

OpenStudy (anonymous):

okayhold on let me see what I get

OpenStudy (anonymous):

(4,1) ?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

yes 4, 1 XD

OpenStudy (anonymous):

thank you!